140702 UCSC Observations

Isaac and Kona arrived at ISB at 1:50pm. The doors to the fog collector had opened but no visible sample was noted. A test run was ran on the collector and it was noted that the Big Fan did not turn on after every test. Upon the third trial the Big Fan turned on. The strand frames and the top plate were removed and cleaned with 409 followed by a heavy rinse with water and DI water. The collector body and the collection jar were sprayed with DI water exclusively. The collector was then reassembled. Further tests were conducted with the Big Fan still having difficulty turning on. The collector was reset though in preparation for the next fog event.

One thought on “140702 UCSC Observations

  1. Kona

    Earlier that morning when Kona had gone in to check on the collector, the system had crashed. After she restarted it and ran a test run it immediately began collecting due to the high humidity that morning. She left it so that she could clean it after the fog was gone.

    Reply

Leave a Reply

Your email address will not be published. Required fields are marked *